# Scaling Laws for Online Advertisement Retrieval

### v1

@@ 4.3. Mapping Function from Model Settings to Computation Cost and FLOPs
1:  Initialize an empty list $results$ to store the results.
@@
10:     Append the result $req_{i}$ to the $results$ list.
@@
12:  return  $results$
5. Applications in Model Designing and Resource Allocation
In this section, we show two real-world applications of the scaling laws identified by the paradigm in section 4 in the advertising system of Kuaishou Technology. To facilitate the description, we first define the following notations:
  • $\mathcal{SP}=[a_{0},a_{1},\ldots,a_{n}]$: A list representing the size parameters of an MLP, where $a_{0}$ denotes the dimension of the MLP input, and $a_{1}$ to $a_{n}$ denote the output dimensions of each layer. Let us denote the TensorFlow meta file generated from the MLP model configured with $\mathcal{SP}$ as $meta_{\mathcal{SP}}$.
  • $F(\mathcal{SP})$: The FLOPs of the network with size parameters $\mathcal{SP}$.
  • $G$: The mapping function from $R/R^{*}$ to the expected online revenue, which is a linear function.
  • $BNSL$: The mapping function from FLOPs to $R/R^{*}$, which follows the form given in Equation 6.
  • $MCET$: The tool (function) used to estimate machine costs, where $MCET(meta_{\mathcal{SP}})$ represents the estimated machine cost for deploying an MLP with size parameters $\mathcal{SP}$.
where $G$, $BNSL$, $F$ and $MCET$ are obtained by referring to section 4.1, section 4.2, and section 4.3, respectively.
5.1. ROI-Constrained Model Designing
For a commercial company, there is usually a standard regarding ROI. Specifically, for an advertising system, we aim for the return on investment (ROI) from the machine to be no less than $\lambda$, where $\lambda$ is typically determined by management based on the company’s operating conditions.
Given a specific MLP model with size parameters $\mathcal{SP}$, the ROI of the model can be estimate by Eq 7:
(7) $\begin{split}ROI&=\frac{G(BNSL(F(\mathcal{SP})))}{MCET(meta_{\mathcal{SP}})}\end{split}$
Thus, given the ROI constraint, we can find the approximate optimal model size by solving the following optimization problem:
(8) $\begin{split}\max_{\mathcal{SP}}&\quad G(BNSL((F(\mathcal{SP}))))\\ \text{s.t. }&\quad\frac{G(BNSL(F(\mathcal{SP})))}{MCET(meta_{\mathcal{SP}})}\geq\lambda\\ \end{split}$
Notice that $\mathcal{SP}$ is a vector where each element represents the number of units in each layer of the MLP. As each element in $\mathcal{SP}$ increases, $F(\mathcal{SP})$, $BNSL(\mathcal{SP})$, and $G(BNSL(\mathcal{SP}))$ also increase, indicating that $F$, $BNSL$, and $G$ are monotonically increasing functions. However, $MCET$ is not a monotonic function and is typically a complex mapping, making the problem non-convex and unsolvable analytically or through binary search.
For the functions in Eq 8, the most time-consuming part is the computation of $MCET(\mathcal{SP})$, which relies on a simulation testing tool. Fortunately, even the computation of $MCET$ does not have a high time and resource cost, allowing us to perform a grid search on this problem. We use 10 T4-GPU machines to run the grid search of about 1000 configurations and find an approximate optimal solution in about two days. Due to space limitations, details of the grid search settings are introduced in appendix B.1.
With the specific ROI constraint of Kuaishou Technology, we solve the approximate optimal size of our foundation MLP model, the $\mathcal{SP}$ of the online base model and the solved model are [3328, 1024, 256, 256, 256, 1] and [16128, 1024, 512, 512, 512, 1], the new model has brought about a 0.85% increase in revenue. Such an improvement is considered significant in our advertising scenario. Without the scaling law, it would be nearly impossible to test the ROI of hundreds of model configurations through online deployment. The time and resource costs of deploying so many models would far outweigh the incremental revenue from the optimized model derived using the scaling law.
5.2. Multi-scenario Resource Allocation
There are often different scenarios that use different models in online advertising systems, such as using different models for different stages and ad placements. Taking different stages as an example, we show the practice of the optimization in multi-scenario resource allocation via scaling law.
Let us assume there are $t$ scenarios, with a total machine cost of $B$. We want to maximize the total revenue through resource allocation across multiple scenarios without increasing the total machine cost, while ensuring that each scenario’s ROI is greater than 1. We can formulate the optimization problem as Eq 9:
(9) $\begin{split}\max_{\{\mathcal{SP}_{i}\mid 0\leq i<t\}}&\sum_{i=0}^{t-1}G(BNSL(F(\mathcal{SP}_{i})))\\ \text{s.t. }&\sum_{i=0}^{t-1}MCET(meta_{\mathcal{SP}_{i}})\leq B\\ \text{s.t. }&\forall i\in\{0,\ldots,t-1\},\frac{G(BNSL(F(\mathcal{SP}_{i})))}{MCET(meta_{\mathcal{SP}_{i}})}\geq 1\\ \end{split}$
This problem can be solved using the same approach as described in sub-section 5.1. The computational complexity of the grid search would be about $t$ times the complexity mentioned in Eq. 8, with the majority of the computation concentrated in the calculation of MCET. In our case, we consider the Matching and Pre-ranking stages as separate scenarios, so $t=2$. Even with $t=2$, we can still use a grid search to solve the problem exhaustively. Due to the initially suboptimal resource allocation, this optimization resulted in a significant 2.8% increase in online revenue. If $t$ is large in practice, making the grid search computationally infeasible, we can employ heuristic information to prune the search space or consider only proportional scaling of the base model. Due to space limitations, more details of multi-scenario resource allocation are shown in appendix B.2.
6. Conclusion
We have systematically studied the scaling laws for online advertisement retrieval and validated the existence of scaling laws in a typical learning-to-rank task with MLP models. To address the challenge of the high cost in time and resources of obtaining scenario-specific scaling law parameters in industrial applications, we propose a lightweight paradigm that transforms most online experiments into low-cost offline experiments through the offline metric R/R* and the machine cost estimation tool. Our work showcases the practical application of the scaling law in optimizing model design under ROI constraints and efficient resource allocation across multiple scenarios in the Kuaishou advertising system, thereby bringing significant commercial value. We also discuss the limitations of our study and outline potential areas for future research in appendix C in the context of scaling laws in online advertisement retrieval.
Acknowledgments
We thank Caiyi Xu, Xiao Liang, Xu He, and Qiang Wang for implementing the machine cost estimation tool and for their daily support of online training and serving issues.
Appendix A More Details about Online Training and Serving
Figure 8 illustrates the pipeline of online training and online serving in the Kuaishou advertising system and the details of ”first-layer optimization” mentioned in section 3.4. The advertising system records logs in real-time for each ad exposure, including the features and labels of the training samples. The downstream training engine reads training data from Kafka storage in real-time and performs online training to optimize the model. Every 20 minutes, the training engine exports the latest model parameters to HDFS for secure storage. Meanwhile, the prediction server periodically polls HDFS to check for new model parameters and updates the local model parameters when a new version is available.
First-layer optimization is an optimization technique designed for MLP models. This technique transforms the computation of the first layer from $FFN(\text{concat}(emb_{user},emb_{ad}))$ (referred to as $FL_{\text{naive}}$) to $FFN(emb_{user})+FFN(emb_{ad})$ (referred to as $FL_{\text{opt}}$). The transformation yields equivalent results but significantly reduces computational overhead. This optimization is applied to the pre-ranking models of the Kuaishou advertising system, where each impression involves evaluating 1500 candidate ads. In this context, for each impression, there is one user embedding ($emb_{user}$) and 1500 ad embeddings ($emb_{ad}$). The operations concat and $+$ are broadcast operations, meaning they can be applied element-wise across arrays of different shapes. Here, the term ”concat” refers to a logical operation with broadcasting properties, which in practice can be implemented using a combination of tf.tile and tf.concat in TensorFlow. By adopting the optimized form $FL_{\text{opt}}$, we can save a substantial amount of computation. To further enhance efficiency, we have developed a scheduled precomputation service called the “Embedding Producer.” This service caches the $FFN(emb_{ad})$ values for all valid ads. It is a CPU-based service that is triggered by model update events and the insertion of new ads. This caching mechanism ensures that the precomputed values are readily available, reducing the computational overhead of the pre-ranking models.
For the training procedure of Pre-ranking models, we have also implemented the first-layer optimization. However, during training, each impression typically involves only about 10 ads, which results in far fewer computational savings compared to the pre-ranking service. For both training and serving, we use T4 GPUs, with each T4 GPU machine equipped with 128 CPU cores. During training, we utilize an asynchronous distributed training framework to improve efficiency and scalability. Both the training and serving frameworks are developed based on TensorFlow. Matching models in our system use the twin-tower structure, inheriting all settings except the first-layer optimization. The inner product of the user and ad tower is directly calculated by brute force for online serving, and no approximate search algorithm such as FAISS is employed.
Figure 8. The pipeline of online training and serving, and details of the first layer optimization for Pre-ranking models.
Appendix B More Details about the Applications of Scaling Law
B.1. ROI-Constrained Model Designing
@@
We set an upper limit on the model FLOPs to 143M and set the step size of the grid search as 16 and 128 for the feature embedding dimensions and the unit number of MLP layers, respectively. This larger step size is chosen because smaller increments have minimal impact on the total FLOPs and result in negligible increases in expected revenue according to the scaling law. The feature embedding dimension ($emb\_dim$) in the grid search starts at 16. The input size of MLP models is determined by multiplying the $emb\_dim$ by the number of sparse features, which is 200 in our experiments, plus an additional 128-dimensional dense feature. For the MLP layers (excluding the output layer, which is fixed at 1), the number of units starts at 128, with the constraint that the number of units in any layer does not exceed the number of units in the previous layer. The size of the first layer outputs does not exceed 1024, because this is the storage limit of the embedding producer. Additionally, we enforce that the number of units in any layer (except the output layer) is at least 1/20 of the number of units in the previous layer as the reason discussed in appendix C.1. We used 10 T4-GPU machines to perform a grid search over approximately 1000 configurations and found an approximate optimal solution in about two days.
B.2. Multi-scenario Resource Allocation
@@
We adopted an MLP model for the Pre-ranking stage and a twin-tower model for the Matching stage. Through offline experiments, we obtained the key functions (including $G$, $BNSL$, etc.) in the scaling laws for both the Matching and Pre-ranking stages using our proposed paradigm. Due to the different model architectures, their machine costs vary in the same size. We used the machine cost estimation tool to additionally measure the machine costs of the twin-tower model at different sizes. Similarly to the estimation of the MLP model, we solved the estimated machine costs of approximately 1000 different sizes of twin-tower models based on the algorithm 1 using 10 T4-GPU machines for about 2 days.
Appendix C Limitations and Future Work
C.1. Optimal Model Designing with Given FLOPs
@@
C.2. New Foundation Models and Joint Laws
While the transformer has become a foundational model in the NLP domain, capable of being applied to almost all tasks and scenarios, the recommendation and advertising fields lack such a foundation model. This absence means that the scaling law parameters obtained in one scenario cannot be readily reused in another. In addition, our scaling laws show that the benefits of increasing FLOPs converge to zero much more rapidly. It’s different from the conclusion in NLP area, where increasing the FLOPs by an order of magnitude typically results in consistent performance improvements. An important reason might be that MLP models in online advertisement retrieval are not as effective foundation models as transformers in NLP. However, it is worth noting that transformers do not universally outperform MLPs in all advertising scenarios, especially when considering the ROI of the machine. Given these challenges, developing more advanced and universal foundation models with better scaling performance (e.g., recent works (Zhai et al., 2024; Zhang et al., 2024b; Guo et al., 2024)) is a crucial direction for future research in online advertisement.
Additionally, obtaining joint laws considering multiple factors such as data, model size, and computational resources is significantly more costly in industrial advertising settings than obtaining single-variable scaling laws. Developing a cost-effective method for deriving joint laws is therefore an important area for future investigation. The high cost of obtaining joint laws is also one of the motivations behind our exploration of unified foundation models.

### v2

@@ D.2 An Detailed Illustration of MCET
1: Initialize an empty list resultsresults to store the results.
@@
10:  Append the result $req_{i}$ to the resultsresults list.
@@
12:return resultsresults
D.3 ROI-Constrained Model Designing
@@
Regrading Case Study 1: We set an upper limit on the model FLOPs to 143M and set the step size of the grid search as 16 and 128 for the feature embedding dimensions and the unit number of MLP layers, respectively. This larger step size is chosen because smaller increments have minimal impact on the total FLOPs and result in negligible increases in expected revenue according to the scaling law. The feature embedding dimension (emb_dimemb\_dim) in the grid search starts at 16. The input size of MLP models is determined by multiplying the emb_dimemb\_dim by the number of sparse features, which is 200 in our experiments, plus an additional 128-dimensional dense feature. For the MLP layers (excluding the output layer, which is fixed at 1), the number of units starts at 128, with the constraint that the number of units in any layer does not exceed the number of units in the previous layer. The size of the first layer outputs does not exceed 1024, because this is the storage limit of the embedding producer. Additionally, we enforce that the number of units in any layer (except the output layer) is at least 1/20 of the number of units in the previous layer, as the reason discussed in appendix E.1. We used 10 T4-GPU machines to perform a grid search over approximately 1000 configurations and found an approximate optimal solution in about two days.
Regrading Case Study 2: We observe from the fitted data points that models in the 0.05B to 0.1B parameter range offer favorable cost-performance trade-offs. Leveraging the scaling law for Transformer-based models, we therefore conduct a grid search within this interval with a step size of 0.01B. Ultimately, a 0.08B-parameter model was deployed. Given that the model size is moderate, we are able to use A10 GPUs for both training and serving. Notably, the scaling law we fit remains in terms of FLOPs and $R/R^{*}$. When the average number of tokens per query is known, there exists a straightforward relationship between model parameter count (in billions) and FLOPs.
D.4 Multi-scenario Resource Allocation
@@
Regrading Case Study 3: As mentioned, we adopted an MLP model for the Pre-ranking stage and a DSSM model for the Matching stage. Through offline experiments, we obtained the key functions (including $G$, BNSLBNSL, etc.) in the scaling laws for both the Matching and Pre-ranking stages using our proposed paradigm. Due to the different model architectures, their machine costs vary in the same size. We used the machine cost estimation tool to additionally measure the machine costs of the DSSM model at different sizes. Similarly to the estimation of the MLP model, we solved the estimated machine costs of approximately 1000 different sizes of DSSM models based on the algorithm 1 using 10 T4-GPU machines for about 2 days.
Regrading Case Study 4: For $Scene_{1}$ and $Scene_{2}$, we perform proportional scaling of the base models in both the Pre-ranking and Matching stages. The scaling ratio ranges from 0.1 to 3.0 with a step size of 0.1. Based on the grid search results and the overall ROI across stages, we ultimately decide to reduce the model size in $Scene_{1}$ by 10% and increase the model size in $Scene_{2}$ by 50%.
Appendix E Limitations and Future Work
E.1 Optimal Model Designing with Given FLOPs
@@
E.2 Jow Laws Identification and Application
Obtaining joint laws considering multiple factors, such as data, model size, and computational resources, is significantly more costly in industrial advertising settings than obtaining single-variable scaling laws. Developing a cost-effective method for deriving joint laws is therefore an important area for future investigation.
An especially promising direction for application is to model the joint law between model size and the computational quota allocation across stages in a cascade ranking system. Such a law could help answer critical questions: How should we rebalance ranking, pre-ranking, and matching stage budgets when scaling up the overall model? At what point does increasing model size in one stage yield diminishing returns due to bottlenecks in another? By formalizing these interactions, joint laws can deepen our understanding of cascade systems and guide principled, globally optimal configuration decisions with only a given budget.
E.3 Beyond Compute Scaling
Although our optimization framework has achieved an overall 5.1% improvement in ad revenue, a substantial portion of this gain arises not from breakthroughs in model architecture, but from correcting historically suboptimal resource allocation and coarse-grained manual tuning. Our results indicate that, unlike in NLP, simply scaling up computation alone is unlikely to bring similar levels of success to the advertising domain for models such as MLP, DSSM, or Transformer.
Therefore, the next major leap in performance will likely require more disruptive innovations. One path is architectural rethinking—designing models with inherently better scaling efficiency, such as sparse, modular architectures [33]. Another is a data-centric paradigm shift: leveraging large language models or product-level innovations to improve the quality, semantic richness, and diversity of training signals. In this view, "scaling" must evolve beyond model size to include the scaling of data intelligence. Only through such foundational changes can we hope to achieve the kind of sustained, huge progress seen in NLP domains.